# Adaptive Canonicalization With Application to Invariant Anisotropic Geometric Networks

### v1

@@ 4.2 Experimental Evaluation of Anisotropic Point Cloud Networks
Table 3: Classification results on ModelNet40.
@@
To evaluate adaptive canonicalization on point cloud classification, we test ModelNet40 [Wu et al., 2015]. The dataset consists of 12,311 shapes from 40 categories, with 9,843 samples for training and 2,468 for testing. We build on two point cloud architectures, PointNet [Qi et al., 2017] and DGCNN [Wang et al., 2019], and apply adaptive canonicalization into their pipeline, denoted respectively AC-PointNet and AC-DGCNN (see Sec. 3.2 and App. E.2). Following Esteves et al. [2018], Deng et al. [2021], we perform on-the-fly rotation augmentation during training, where the dataset size remains unchanged, and for the test set, each test example is arbitrarily rotated. We compare with PointNet, DGCNN, equivariant networks VN-PointNet and VN-DGCNN from Deng et al. [2021], canonicalization methods CN-PointNet and CN-DGCNN from Kaba et al. [2023], and traditional augmentation baselines where the training set is statically expanded with pre-generated rotated samples, denoted PointNet-Aug and DGCNN-Aug. For further details, see App. F. Tab. 3 shows the classification performance. We see our method outperforms both equivariant models and canonicalization baselines, as it allows the model to learn both local geometric features and optimal reference alignments for each class.
@@ 5 Conclusions
Limitations and Future Work. Our framework is naturally suited to classification tasks, and at the current scope of the paper, we did not address regression tasks. We will extend the adaptive canonicalization to regression in future work. Another limitation of our approach is that prior maximization requires solving $D$ optimizations at runtime for the $D$ classes. In future work, we will reduce this to a single optimization to improve efficiency.
@@ Acknowledgments
This supplemental material is structured as follows:
  • Appendix A provides an extended review of related work in equivariant machine learning, including further discussions on canonicalization methods, frame-averaging, and equivariant architectures.
  • Appendix B presents additional theoretical analysis of universal approximation theorems. We discuss the universal approximation of Euclidean functions as well as multi-sets functions.
  • Appendix C includes additional examples of continuous prior maximization. Here, we present the continuous prior maximization for unbounded point clouds under rotations, for continuous to discrete images with rotations and other transformations, as well as for discrete to discrete images with rotations and other image transformations.
  • Appendix D reviews the functional calculus and demonstrates how to plug self-adjoint operators inside functions, specifically in the context of the spectral filters. In addition, we also include a discussion on the automorphism invariance of the spectral filter.
  • Appendix E describes the construction of adaptive canonicalization applied to spectral graph neural networks and point cloud networks, as discussed in Sec. 3. We introduce how to construct the anisotropic nonlinear spectral filters and how the adaptive canonicalization method is applied to PointNet and DGCNN.
  • Appendix F outlines the experimental setup and details for the experiments in Sec. 4. We report the used datasets, including the toy problem, TUDataset, and OGB benchmarks. We include the dataset split, competing methods, and hyperparameters.
  • Appendix G presents additional results of our method, including trade-offs between sampling- and optimization-based adaptive canonicalization, the node-level anisotropic spectral filters, and the generalization of point cloud under out-of-distribution rotations.
@@ Appendix A Related Work
A.1 Canonicalization Methods: Fixed and Learned Approaches
@@
Canonicalization [Babai and Luks, 1983, Palais and Terng, 1987, Ma et al., 2023, 2024] is a classical strategy for handling symmetry in data, especially for tasks where invariance or equivariance to group actions is desirable [Gerken et al., 2023]. It preprocesses each input by mapping it to a standard form prior to downstream learning and inference, so that all symmetry-equivalent inputs are treated identically by the subsequent model. There are two common ways for canonicalization: fixed or learned approaches.
@@
  • Fixed Canonicalization. Fixed canonicalization uses deterministic and often analytic procedures to assign a unique representative to each symmetry orbit. For example, principal component analysis (PCA) [Jolliffe and Cadima, 2016] alignment canonically orients an object (e.g., a point cloud or molecule) by rotating it so its principal components align with the coordinate axes [Kazhdan et al., 2003]. Procrustes analysis [Gower, 1975] canonically orients sets of points by finding the optimal rotation, translation, and scale that minimizes squared point-to-point distances to a reference. For sets and graphs, canonicalization can be achieved by reordering nodes, atoms, or features so that isomorphic inputs share a single labeling. In spectral methods and spectral graph neural networks where eigenvectors are fundamental [Kipf, 2016, Defferrard et al., 2016a, Von Luxburg, 2007, Belkin and Niyogi, 2003, Dwivedi et al., 2023, Maskey et al., 2022], canonicalization of spectral decomposition [Lim et al., 2022, 2023, Ma et al., 2023, 2024] addresses eigenbasis ambiguity [Chung, 1997, Spielman, 2012] by processing each eigenspace independently and selecting representative eigenvectors or directions by applying orthogonal or axis-based projections, typically as a graph preprocessing step. An alternative approach is eigenbasis canonicalization via the input signal, where the signal itself is used to define a canonical spectral representation, making the spectral transformation independent of the arbitrary choice of eigenvectors [Lin et al., 2024a, Geisler et al., 2024].
@@
  • Learned Canonicalization. Learned canonicalization [Zhang et al., 2018, Kaba et al., 2023, Luo et al., 2022] seeks to overcome the rigidity and inflexibility of fixed rules with a trainable mapping that selects a representative for each symmetry orbit. The canonicalizer is parameterized (typically as a neural network and trained to produce canonical forms. For example, Kaba et al. [2023] developed a neural network that learns the canonicalization transformation, which enables plug-and-play equivariance, e.g., orthogonalizing learned features via the Gram-Schmidt process [Trefethen and Bau, 2022]. Their results show that the learned canonicalizers outperform fixed canonicalizers.
@@
However, Dym et al. [2024] pointed out that regardless of whether the canonicalizations are learned or not, a continuous canonicalization does not exist for many common groups (e.g., Sn, SO(d), O(d) on point clouds $n\geq d$). Therefore, while learned canonicalization improves empirical performance, it remains generally discontinuous and can induce instability, hinder generalization, limit model reliability on boundary cases, or out-of-distribution data. In contrast, our adaptive canonicalization framework learns the optimal transformation for each input by maximizing the predictive confidence of the network, resulting in a continuous and symmetry preserving mapping.
Symmetry Adaptation for Pretrained Models via Canonicalization.
@@
The recent work [Mondal et al., 2023] addresses symmetry adaptation in large-scale networks by learned canonicalization before input is processed by a pretrained network. Therefore, it can convert pretrained deep networks to handle symmetry, bypassing the need to redesign or retrain the network. However, the canonicalization mapping itself is not continuous, and misalignments between canonical forms and the pretrained model’s priors can still limit effectiveness. In addition, it is important to note that its canonical form is determined by the canonicalizer’s internal objective and (potentially) supplemented with dataset priors. In contrast, our adaptive canonicalization based on prior maximization adapts the canonicalization to the downstream predictor by explicitly selecting, for each input, the transformation that maximizes the predictive confidence of the prediction model for inference. Another recent extension of canonization work [Singhal et al., 2025] explores a set of transformations at test time and uses the scoring functions of large pre-trained foundation models like CLIP [Radford et al., 2021] or SAM [Kirillov et al., 2023] to select the most “canonical” representation upon which downstream inference is performed. Note that the work in Singhal et al. [2025] does not involve any network retraining, uses the foundation models as is, and performs canonicalization entirely at inference by optimizing over transformations. While it achieves strong empirical performance, its canonicalization mapping is not continuous, and therefore, small input changes could cause abrupt switches in the selected canonical view.
@@ A.2 Frame Averaging
Frame averaging [Puny et al., 2021] achieves equivariance to group symmetries by averaging a network’s output over a set of group transformations (known as a “frame”). It is built on the classical group averaging operator, which guarantees symmetries by summing a function over all group elements. Frame averaging has two main advantages: 1) it allows adaptation of standard non-equivalent network architectures to handle symmetry, similar to canonicalization methods, and 2) it avoids computational intractability of full group averaging, especially for large or continuous groups. Recent work [Lin et al., 2024b] proposes minimal frame averaging that attains strong symmetry coverage with small frames. Domain-specific frame averaging methods [Duval et al., 2023b, Atzmon et al., 2022] show that it can be deployed in material modeling and geometric shape analysis. However, it requires a careful selection of a suitable frame. In addition, frame averaging uses a fixed set of transformations independent of the input or task, potentially leading to sub-optimal or less discriminative feature representations. In contrast, our adaptive canonicalization learns the optimal transformation for each input in a data- and network-dependent way, yielding symmetry preserving continuous functions that can improve representation quality and empirical task performance.
@@ A.3 Equivariant Architectures
Equivariant architectures are a class of models explicitly designed to respect symmetry groups acting on the data. Formally, a network is equivariant to a group of transformations if, when the input is transformed by a symmetry group action, the output transforms via the same group action. That is, for a group $G$ and a function $f$, equivariance guarantees $f(g\cdot x)=g\cdot f(x)$ for all $g\in G$ and input $x$. It has been developed across images [Cohen and Welling, 2016a, b, Kondor and Trivedi, 2018, Worrall et al., 2017, Weiler et al., 2018b], graphs [Bronstein et al., 2017, Zaheer et al., 2017, Gilmer et al., 2017, Maron et al., 2019, Kofinas et al., 2024, Thiede et al., 2020, Vignac et al., 2020, Keriven and Peyré, 2019], molecules [Thomas et al., 2018, Brandstetter et al., 2021, Anderson et al., 2019, Fuchs et al., 2020, Satorras et al., 2021, Duval et al., 2023a, Schütt et al., 2017, Liao and Smidt, 2023, Hordan et al., 2025, Du et al., 2022, Passaro and Zitnick, 2023], and manifolds [Masci et al., 2015, Monti et al., 2017, Cohen et al., 2019, Weiler et al., 2021]. Notably, equivariant networks often demand group-specific designs that rely on group theory, representation theory, and tensor algebra. This can reduce flexibility and raise compute and memory costs [Pertigkiozoglou et al., 2024, Wang et al., 2023, Liao and Smidt, 2023]. Making nonlinearities, pooling, and attention strictly equivariant further constrains layer choices and can increase parameter count and runtime. Moreover, imposing symmetry throughout the stack may limit the expressivity when data only approximately respect the assumed symmetry or contain symmetry-breaking noise [Wang et al., 2022, Lawrence et al., 2025]. In contrast, our approach handles symmetry by learning an input- and task-dependent canonical form through prior maximization and applying a standard backbone. By construction, this mapping is continuous and symmetry preserving. Our adaptive canonicalization removes heavy group-specific layers, reduces per-layer equivariant cost, and keeps ordinary nonlinearities and pooling.

### v2

@@ 5.2 Experimental Evaluation of Anisotropic Point Cloud Networks
Table 3: Classification results on ModelNet40. Results of competing methods marked with * are taken from deng2021vector, luo2022equivariant, kaba2023equivariance.
@@
[…]
To evaluate adaptive canonicalization on point cloud classification, we test ModelNet40 [wu20153d]. The dataset consists of 12,311 shapes from 40 categories, with 9,843 samples for training and 2,468 for testing. We build on two point cloud architectures, PointNet [qi2017pointnet] and DGCNN [wang2019dynamic], and apply adaptive canonicalization into their pipeline, denoted respectively AC-PointNet and AC-DGCNN (see Sec. 4.2 and App. E.4). Following esteves2018learning, deng2021vector, we perform on-the-fly rotation augmentation during training, where the dataset size remains unchanged, and for the test set, each test example is arbitrarily rotated. We compare with PointNet, DGCNN, equivariant networks VN-PointNet and VN-DGCNN from deng2021vector, canonicalization methods CN-PointNet and CN-DGCNN from kaba2023equivariance, and traditional augmentation baselines where the training set is statically expanded with pre-generated rotated samples, denoted PointNet-Aug and DGCNN-Aug. For further details, see App. F. Tab. 3 shows the classification performance. We observe that our method outperforms the competing baselines and it allows the model to learn both local geometric features and optimal reference alignments for each class.
@@ 6 Conclusions
Limitations and Future Work.
Our framework is naturally suited to classification tasks, and at the current scope of the paper, we did not address regression tasks. We will extend the adaptive canonicalization to regression in future work. Another limitation of our approach is that prior maximization requires solving $D$ optimizations at runtime for the $D$ classes. In future work, we will reduce this to a single optimization to improve efficiency.
@@ Appendix A Related Work
A.1 Canonicalization
@@
Canonicalization [babai1983canonical, palais1987general, ma2023laplacian, ma2024canonicalization] is a classical strategy for handling symmetry in data, especially for tasks where invariance or equivariance to group actions is desirable [gerken2023geometric]. It preprocesses each input by mapping it to a standard form prior to downstream learning and inference, so that all symmetry-equivalent inputs are treated identically by the subsequent model. There are two common ways for canonicalization: fixed or learned approaches.
@@
  • Fixed Canonicalization. Fixed canonicalization uses deterministic and often analytic procedures to assign a unique representative to each symmetry orbit. For example, principal component analysis (PCA) [jolliffe2016principal] alignment canonically orients an object (e.g., a point cloud or molecule) by rotating it so its principal components align with the coordinate axes [kazhdan2003rotation]. Procrustes analysis [gower1975generalized] canonically orients sets of points by finding the optimal rotation, translation, and scale that minimizes squared point-to-point distances to a reference. For sets and graphs, canonicalization can be achieved by reordering nodes, atoms, or features so that isomorphic inputs share a single labeling. In spectral methods and spectral graph neural networks where eigenvectors are fundamental [kipf2016semi, defferrard2016convolutional, von2007tutorial, belkin2003laplacian, dwivedi2023benchmarking, maskey2022generalized], canonicalization of spectral decomposition [lim2022sign, lim2023expressive, ma2023laplacian, ma2024canonicalization] addresses eigenbasis ambiguity [chung1997spectral, spielman2012spectral] by processing each eigenspace independently and selecting representative eigenvectors or directions by applying orthogonal or axis-based projections, typically as a graph preprocessing step. An alternative approach is eigenbasis canonicalization via the input signal, where the signal itself is used to define a canonical spectral representation, making the spectral transformation independent of the arbitrary choice of eigenvectors [lin2024equivariant, geisler2024spatio].
@@
  • Learned Canonicalization. Learned canonicalization [zhang2018learning, kaba2023equivariance, luo2022equivariant] seeks to overcome the rigidity and inflexibility of fixed rules with a trainable mapping that selects a representative for each symmetry orbit. The canonicalizer is parameterized (typically as a neural network and trained to produce canonical forms. For example, kaba2023equivariance developed a neural network that learns the canonicalization transformation, which enables plug-and-play equivariance, e.g., orthogonalizing learned features via the Gram-Schmidt process [trefethen2022numerical]. Their results show that the learned canonicalizers outperform fixed canonicalizers.
@@
However, dym2024equivariant pointed out that regardless of whether the canonicalizations are learned or not, a continuous canonicalization does not exist for many common groups (e.g. Sn, SO(d), O(d) on point clouds $n\geq d$). Therefore, while learned canonicalization improves empirical performance, it remains generally discontinuous and can induce instability, hinder generalization, limit model reliability on boundary cases, or out-of-distribution data. In contrast, our adaptive canonicalization framework learns the optimal transformation for each input by maximizing the predictive confidence of the network, resulting in a continuous and symmetry preserving mapping. We include a detailed comparison of most related canonicalization work with our method below.
A.1.1 Equivariance with Learned Canonicalization Functions
Recently, kaba2023equivariance introduced energy-based canonicalization. The central idea is to learn an energy function over samples and group elements, and define the canonicalization as minimizing this energy with respect to the group, given a fixed datapoint. Specifically, their energy minimization is related to our prior maximization adaptive canonicalization method. However, there are several important conceptual and technical differences between their approach and our adaptive canonicalization.
First, their energy $s$ is not the task neural network like in our analysis, but rather some other trainable neural network. Similarly to our approach, training $s$ end-to-end with the task neural network can be seen as canonicalization that depends on the task network, if one considers the full end-to-end architecture consisting both of the energy minimization and the task network. However, the approach in kaba2023equivariance does not give continuity guarantees as opposed to our approach. Second, in their work, they consider symmetries based on group actions, while we consider a more general setting of canonicalization transformations (or augmentations) that need not be based on groups. This makes our approach much more applicable across different domains. Moreover, in their framework, there is one canonical form for each datapoint and network, while in our approach, each output channel of the network defines a different canonical form of the datapoint. This allows our approach to preserve continuity. Notably, kaba2023equivariance do not attempt to study the continuity of the end-to-end predictor.
Another difference in kaba2023equivariance is that when training is initialized, the canonicalizing energy $s$ is random. This leads each datapoint to be randomly transformed, so the task neural network initially has to perform well at all orientations of the data. This can lead the task network to ultimately learn an “average behavior,” not specializing in any special orientation but rather performing reasonably well on all orientations of the data. In other words, the limited set of trainable parameters has to simultaneously specialize in many orientations, which reduces the network’s expressive power. In contrast, in our prior maximization approach, from the beginning of training, the network only pays a price for not performing well on the single best orientation per datapoint (on which the network performs the best). This encourages the network to specialize on one canonical orientation per datapoint, and not learn an average behavior. Hence, in our approach, all trainable parameters of the task network can focus on performing well only on the sole canonical orientation of each datapoint.
A.1.2 Canonicalization and Data Re-Alignments
As discussed in App. A.1.1, in the energy-based canonicalization framework of [kaba2023equivariance], the canonicalizing energy $s$ is random at initialization. As a result, it leads each datapoint to be randomly transformed and the task neural network initially has to perform well at all orientations. This can lead the task network to ultimately learn an “average behavior,” not specializing in any special orientation but rather performing reasonably well on all orientations of the data. To address this effect, mondal2023equivariant biases the canonical transformation of each datapoint to be the identity, assuming that the datapoints in the training set already have a small orientation variance. On top of that, schmidt2025robust iteratively reduces the orientation variance of the training set by iteratively reorienting datapoints that lead to a large loss. We note that these approaches are rather different from our prior maximization method, and they do not try to address the continuity problem in canonicalization.
A.1.3 Weighted Canonicalization
The energy-based canonicalization [kaba2023equivariance] was further explored and extended in [shumaylov2024lie] on symmetries defined by general Lie group actions. Similarly to our work, shumaylov2024lie also discusses continuity preservation, but their approach is different from ours. In the work of shumaylov2024lie, they define the notion of weighted canonicalization, which is a similar concept to the weighted frame introduced by dym2024equivariant. Here, to each datapoint there is an assigned probability measure over the orbit of the datapoint. Namely, the distribution is over the space of data instead of over the group like in the weighted frames of dym2024equivariant. With respect to energy minimization, this approach is not very different from kaba2023equivariance. The main difference is in the minimization algorithm, which minimizes over the Lie algebra instead of the Lie group. It is important to note that their work does not train the energy end-to-end with the neural network. Hence, in their work, the canonicalization depends on the whole training set, but not on the task network, which is quite different from our approach. Moreover, in their setup, one has to learn an approximation of the data distribution, which is invariant to the group action. This is a highly nontrivial approach to implement. In contrast, our prior maximization is simple and direct.
A.1.4 Test-Time Canonicalization
@@
Another recent extension of canonicalization, [singhal2025test] , explores a set of transformations at test time and uses the scoring functions of large pre-trained foundation models like CLIP [radford2021learning] or SAM [kirillov2023segment] to select the most “canonical” representation upon which downstream inference is performed. Note that the work in singhal2025test does not involve network retraining, uses the foundation models as is, and performs canonicalization entirely at inference by optimizing over transformations. While it achieves strong empirical performance, its canonicalization mapping is not guaranteed to be continuous, and in fact, continuity is not discussed. Therefore, small input changes may cause abrupt switches in the selected canonical view.
A related line of work is inverse transformation search [schmidt2024tilt], which also performs test-time optimization over transformations to exploit invariances. Their method focuses on the standard action of the special linear group on images (rotations, scalings, and shear transformations), i.e., purely group-based symmetries, and similar to singhal2025test, does not address continuity. In addition, their method does not train the model simultaneously with the canonicalization. While their energy-induced confidence is similar to our prior-maximization formulation in the classification setting, it does not lead to a continuity guarantee. In contrast, our work rigorously develops sufficient conditions on the canonicalizer for the canonicalized network to be continuous and have a universal approximation property. Specifically, our one-vs.-all setting is a different type of energy that does lead to continuous end-to-end classifiers.
@@ A.2 Frame Averaging
Frame averaging [puny2021frame] achieves equivariance to group symmetries by averaging a network’s output over a set of group transformations (known as a “frame”). It is built on the classical group averaging operator, which guarantees symmetries by summing a function over all group elements. Frame averaging has two main advantages: 1) it allows adaptation of standard non-equivalent network architectures to handle symmetry, similar to canonicalization methods, and 2) it avoids computational intractability of full group averaging, especially for large or continuous groups. Recent work [lin2024equivariance] proposes minimal frame averaging that attains strong symmetry coverage with small frames. Domain-specific frame averaging methods [duval2023faenet, atzmon2022frame] show that it can be deployed in material modeling and geometric shape analysis. However, it requires a careful selection of a suitable frame. In addition, frame averaging uses a fixed set of transformations independent of the input or task, potentially leading to sub-optimal or less discriminative feature representations. In contrast, our adaptive canonicalization learns the optimal transformation for each input in a data- and network-dependent way, yielding symmetry preserving continuous functions that can improve representation quality and empirical task performance.
A related work that addresses continuity is the weighted frame averaging proposed by dym2024equivariant. They first prove that in many well-known cases, continuous canonicalization is impossible. This does not contradict our work, as in dym2024equivariant the canonicalization is a function solely of the datapoint, and not the task network. Then, they define a variant of frame averaging, called weighted frame averaging, in which to each datapoint there is an associated probability distribution over the group, and the frame averaging is performed with respect to this measure. This construction yields continuity guarantees. However, their focus is fundamentally different from ours: they study frame averaging while we focus on canonicalization. Moreover, the weighted frame averaging is a function only of the data, not the neural network, as opposed to our method. We next compare our framework to weighted frame averaging in more detail.
  • In our method, data need not come from a vector space of a fixed dimension (see e.g., the application for graphs). In contrast, weighted frame averaging requires working with data that comes from a vector space of a fixed dimension.
  • In our work, symmetries need not be based on group representations. Our notion of symmetry is called a transformation family, and our “symmetries” are not even required to be invertible or based on a group action. See, for example, the image truncation transformation in App. E.5. On the other hand, the symmetries in weighted frame averaging are required to be representations of compact groups.
  • In the framework of weighted frame averaging, the requirement that the weighted frame is robust is sufficient for continuity of the canonicalized function. However, this requirement is quite strong, and constructing robust frames could be challenging. In contrast, we achieve continuity of the canonicalized function even though the maximizer in prior maximization need not be continuous in any sense (as a function of the datapoint). The maximizer need not even be uniquely defined. Hence, practitioners can use our method out of the box on new domains with new symmetries without having to prove any nontrivial mathematical results. In weighted frame averaging, employing the method in a new setting typically requires carefully constructing a problem-specific weighted frame and proving that it is robust. It often involves nontrivial mathematical work and there is no general recipe for doing so. In our case, once an architecture and a symmetry setting are chosen, prior-maximization adaptive canonicalization is straightforward to implement and does not require additional sophisticated proofs from the practitioner. The only assumption that needs to be checked is continuity of the chosen transformations, which is usually easy to verify.
  • Robust weighted frames often require averaging the predicting network over many transformations of the input. In practice, our prior maximization approach works with fewer random argmax candidates. For comparison, frame averaging for rotations of $n$ points in a 3D point cloud requires an order of $n^{2}$ transformations (where $n=1024$ in the ModelNet40 dataset for example), while, in practice, prior maximization performs well with a total of 50 random transformations.
These differences imply that our framework is more flexible in terms of the data types and symmetry structures it can accommodate, while also imposing a lighter mathematical burden on practitioners who wish to apply it. This gives much more freedom for future practitioners, and may lead to a wider adaptation of the method.
We note that the earlier work [basu2023efficient] proposed a similar idea to weighted frame averaging, but did not study continuity preservation.
@@ A.3 Equivariant Architectures
Equivariant architectures are a class of models explicitly designed to respect symmetry groups acting on the data. Formally, a network is equivariant to a group of transformations if, when the input is transformed by a symmetry group action, the output transforms via the same group action. That is, for a group $G$ and a function $f$, equivariance guarantees $f(g\cdot x)=g\cdot f(x)$ for all $g\in G$ and input $x$. It has been developed across images [cohen2016group, cohen2016steerable, kondor2018generalization, worrall2017harmonic, weiler2018learning], graphs [bronstein2017geometric, zaheer2017deep, gilmer2017neural, maron2018invariant, kofinas2024graph, thiede2020general, vignac2020building, keriven2019universal], molecules [thomas2018tensor, brandstetter2021geometric, anderson2019cormorant, fuchs2020se, satorras2021n, duval2023hitchhiker, schutt2017quantum, liao2023equiformer, hordan2025spectral, du2022se, passaro2023reducing], and manifolds [masci2015geodesic, monti2017geometric, cohen2019gauge, weiler2021coordinate]. Notably, maron2019universality studies the universal approximation property for equivariant architectures. Notably, equivariant networks often demand group-specific designs that rely on group theory, representation theory, and tensor algebra. This can reduce flexibility and raise compute and memory costs [pertigkiozoglou2024improving, wang2023discovering, liao2023equiformer]. Making nonlinearities, pooling, and attention strictly equivariant further constrains layer choices and can increase parameter count and runtime. Moreover, imposing symmetry throughout the stack may limit the expressivity when data only approximately respect the assumed symmetry or contain symmetry-breaking noise [wang2022approximately, lawrence2025improving]. In contrast, our approach handles symmetry by learning an input- and task-dependent canonical form through prior maximization and apply a standard backbone. By construction, this mapping is continuous and symmetry preserving. Our adaptive canonicalization removes heavy group-specific layers, reduces per-layer equivariant cost, and keeps ordinary nonlinearities and pooling.